# Non-normal Dynamics on Non-reciprocal Networks: Reactivity and Effective Dimensionality in Neural Circuits

### v1

@@ III.3 Reactivity
Reactivity is one of the distinctive markers of non-normality since it only arises in its presence. However, the reverse is not always true, i.e. not all non-normal systems exhibit reactivity. A reactive system is one in which, during a transient phase the system moves away from the steady state, following a perturbation. More specifically, at small time scales the 222-norm of a variable is enhanced before exponentially decaying back to the equilibrium point. Formally, reactivity occurs only when the maximum eigenvalue of the Hermitian part of the Jacobian, known as numerical abscissa $m$, is positive:
$m(\mathbf{J})=\sup\sigma\left[\left(\mathbf{J}+\mathbf{J}^{*}\right)/2\right],$ (19)
where $\sigma(\mathbf{J})$ denotes the spectrum of the matrix $\mathbf{J}$ and $\mathbf{J}^{*}$ is its conjugate transpose. The growth during the transient is proportional to the magnitude of $m(\mathbf{J})$ (as long as $m(\mathbf{J})>0$).
In the cyclic network, the numerical abscissa can be evaluated analytically. In particular, we find that for the relevant portion of parameter space $\gamma_{\mu},\gamma_{\nu}>0$, we obtain:
@@
The results are summarized in Fig. 3, where we present the level curves of $m$ for several values of $\gamma_{\ell}$ and for $c=0,\epsilon$. Notably, panel f) indicates that the behavior of the feed-forward system as a function of $m$ –whose analytic calculation is much more involved– qualitatively matches with that of the cyclic case, shown in panel e). This observation suggests that Eq.20 may serve as a qualitative reference even for intermediate configurations with $0<c<\epsilon$.
@@
Interestingly, it turns out that reactivity is governed by two distinct quantities: $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, with the absolute value of their ratio being the non-normality parameter. This implies that, although non-normality is needed to trigger reactivity, a fixed non-normality level does not correspond to a single level of reactivity. In other words, the ratio of $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$ can remain constant while adjusting both numerator and denominator to enhance $m$. This is shown explicitly in Fig. 3 panels e) and f), where the lines representing constant values of $\xi$ or $\zeta$ cross the level curves of $m$, i.e., while moving along a line with constant non-normality, multiple values of reactivity can be achieved.
@@
Moreover, from Eq. 20 we define $\Gamma\equiv\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, i.e. the sum of all coupling constants. In what follows we fix the non-normality and distinguish between two regimes: one with weak interactions and weak reactivity ($\Gamma,m\ll 1$, illustrated in Fig. 3 panels c), e)) and one of strong interactions and strong reactivity ($\Gamma,m\gg 1$, illustrated in panels d), f)).
@@
In the right dashed region, we show the time-evolution of the total activity $\rho(t)$ $=\sqrt{\sum_{i}x_{i}^{2}+y^{2}_{i}}$, starting from an initial condition close to the bisector, where Eq. 20 holds. In the first row, parameters are chosen so that the inner systems are all equivalent, and in particular, the non-normality parameter of the inner systems, $\zeta$, is kept constant. We want to compare the evolution of the norm $\rho(t)$ when the reactivity is small (panel c)) with a case in which reactivity is large (panel d)). As expected we observe that, when the reactivity is large, the system moves away from the fixed point in 0 before relaxing. From these plots, the cyclic system seems to be more reactive than the decoupled system, and that reactivity increases further as $c$ approaches $\epsilon$.
However, the reactivity may appear larger for the coupled systems, simply because the coupling itself increases the non-normality, which has an effect in the reactivity. To isolate the effects of reactivity, in the lower row, we fix the non-normality of each system to the same value. In particular we choose parameters in the cyclic system such that $\xi=\zeta$, i.e. we adjust $\gamma_{\mu}$ to compensate for the introduction of $\gamma_{\ell}$.
Again, the left panel, panel e) is for small values of reactivity, and the right panel, panel f) is for larger values of the reactivity, showing larger excursions away from the fixed point. Two main messages emerge from these plots: first, we verify that the cyclic system can be reduced to the decoupled case merely through coupling parameter shifts; second, we show that the feedforward system undergoes a genuine increase in reactivity when internal non-normality is held constant.
For the feedforward system, instead, we do not have a unique global parameter to quantify the total non-normality and summing the parameters for inner and outer non-normality would be an arbitrary choice. A specific choice of the couplings that makes the reactivity coincide surely exists, but estimating it is not easy. The point that we want to make here is that the external source of non-normality enhances the reactivity.
In conclusion, the feed-forward network actually increases reactivity compared to the decoupled system. In contrast, in the cyclic system the enhancement of reactivity can be fully accounted for by equalizing the level of non-normality.
@@ III.4 Analysis of the stochastic evolution of the dynamics
As shown in [13, 7], the microscopic underlying system shows very strong finite-size effects. In particular it was shown in [7] using Eq. 3, that the demographic noise induces a minimum of the effective potential is the origin, thus generating a bistability in the system. In presence of strong non-normality, the shear flow along the diagonal, associated with the two nearly-collapsing eigenvectors, facilitates the jumps between the two fixed points and the dynamics resembles the avalanching behaviors observed in neural system, where the global activity undergoes large irregular fluctuations between a low-activity and a high-activity state.
Here we want to explore the behavior of the coupled system in response to noise. We then couple three mean-field dynamics, where each obeys the system size expansion as in [13]:
@@
$\frac{dy_{i}}{dt}$ $=-\alpha y_{i}+(1-y_{i})f(s_{i})+\sqrt{\alpha y_{i}+(1-y_{i})f(s_{i})}\eta_{y_{i}}.$
Also in this case, the introduction of noise induces a new stable equilibrium point in zero for all parameter choices, given that $h$ is kept small. Again, when the deterministic stable point bifurcates from zero, bistability appears. Bistability is not a sufficient ingredient to observe avalanching behavior: a high non-normality is needed [7]. However, for too high values of $\xi$ and $\zeta$ the system tends to be trapped in the active minimum. Since we are interested in how the activity propagates across populations, we used the fraction of active excitatory neurons to visualize the trajectories, Fig. 4. Thanks to the non-normality, similar plots could be obtained for inhibitory activities.
@@
Once a sufficiently large non-normality level is set, we can distinguish between two separate behaviours, one where the trajectories are incoherent and one where the system shows dynamical patterns, as a function of the parameter $\gamma_{\ell}$. When $\gamma_{\ell}<<\gamma_{\mu},\gamma_{\nu}$ the system lies in an incoherent phase (Fig. 4 panels b) and d)), while when $\gamma_{\ell}\sim\gamma_{\mu},\gamma_{\nu}$ the behavior is more coherent (Fig. 4 panels a) and c)). When the connectivity among units is large enough and the architecture is cyclic, the activation between units is highly correlated (Fig. 4a)). Instead, in the feed-forward case, when the connectivity among units is large enough, the units are activated sequentially. To fix the ideas, let us take two adjacent populations, where the bottom one projects input to –but does not receive any input from– the top one. When the bottom population transits to the active minimum, it activates also the upper population. The process propagates along the chain, giving rise to the step-like trajectories we observe in Fig. 4c). A discussion about the effective potential induced by the noise and the external couplings is detailed in SM.
Figure 4: Comparison of cyclic and feedforward setups in the stochastic regime. The top panels illustrate the cyclic configuration, while the bottom ones refer to the feedforward setup. The comparison between the left and the right columns shows how the system can be driven from an incoherent regime (on the right, panels b) and d)) to a coherent phase in THE cyclic setup (panel a)) and to step-like trajectories in the feed-forward one (panel c)). Parameter choices are as follows: a) $\gamma_{\ell}=1.13$ and b) $\gamma_{\ell}=0.13$; for both $\xi=0.004$. c) $\gamma_{\ell}=1.13$ and d) $\gamma_{\ell}=0.07$; for both $\zeta=0.005$.
@@ IV Discussion and conclusions
Firstly, through the analysis of bifurcations and reactivity, we find that systems with feedback are more controllable, meaning that changes in parameter settings – or changes in physiological conditions – result in smoother alterations of network responses compared to purely feedforward systems. This enhanced controllability implies that the activation patterns of neurons –which are the substrate for neural representation of the stimulus– are more robust to perturbations when feedback is present. Similarly, it has been proposed that when stimuli become consciously reportable, they elicit an ‘ignition process’, which causes information about a brief stimulus to become sustained and broadcasted across many brain areas [30]. Our work suggests, in an extremely simplified setup, that the presence of feedback might be a key ingredient in such process. More specifically, our model reveals that the introduction of feedback into the system plays a critical role in generating coherence between processing stages. This observation aligns with experimental findings of ‘zero-lag interaction’ between layers during perception and visual working memory [31, 32]. These are frequently attributed to either common external inputs or specific network motifs [31, 33]. Interestingly, our model suggests that feedback alone might suffice to produce this synchrony, even in the absence of external modulation. Exploring the influence of time-lagged feedback would be a valuable future direction to test the validity of this hypothesis.
Additionally, our model shows that in a purely feedforward setup, the stochastic dynamics exhibits a greater degree of variability and complexity in the system’s response patterns, compared to the cyclic setup. In the absence of feedback, the neural activity is more distributed across a broader range of activation patterns, reflecting a more diverse set of responses to sensory inputs.
Remarkably, when the inter-unit connectivity is weak, coherent representations are completely lost, both for feedforward and cyclic networks.
A priori, we chose $n=3$ to investigate the potential role of complex eigenvalues and to explore whether the system could sustain quasi-cycles under the influence of noise. However, our analysis reveals that, despite the theoretical presence of complex eigenvalues, no oscillatory behavior emerges in practice. As $n$ increases, the frequency of the oscillation decreases, and the quasi-cycle dynamics potentially predicted by the mathematical formalism might be observable for larger $n$. This could be investigated with numerical simulations of systems with $n>3$, but is beyond the scope of the present work.
This study contributes to the growing body of literature suggesting that feedback plays a pivotal role in the neural dynamics underlying perception. Our model provides new insights into how feedback mechanisms could inherently promote synchrony and robustness in neural networks, even in the absence of common input.

### v2

@@ III.3 Reactivity
To quantify the immediate transient amplification capability of the network, we calculate the reactivity $\mathcal{R}$, defined as the maximum eigenvalue of the Hermitian part of the Jacobian [12]:
$\mathcal{R}=\lambda_{\max}\left(\mathcal{H}(J)\right)=\lambda_{\max}\left(\frac{J+J^{\top}}{2}\right).$ (21)
While the spectral abscissa $s(J)=\max_{i}\text{Re}(\lambda_{i})$ determines the asymptotic stability ($t\to\infty$), the reactivity governs the instantaneous growth rate of the perturbation magnitude at $t=0$. This is explicit in the time evolution of the Euclidean norm $\|\delta\mathbf{x}\|^{2}$:
$\frac{1}{2}\frac{d}{dt}\|\delta\mathbf{x}\|^{2}=\delta\mathbf{x}^{\top}\left(\frac{J+J^{\top}}{2}\right)\delta\mathbf{x}\leq\mathcal{R}\|\delta\mathbf{x}\|^{2}.$ (22)
A positive reactivity $\mathcal{R}>0$ in an asymptotically stable system ($s(J)<0$) implies a non-monotonic relaxation to the steady state. In the language of critical phenomena, this corresponds to a regime of transient fluctuation amplification: even though the system is formally in the disordered (stable) phase, the violation of detailed balance drives a short-time divergence of the susceptibility, allowing fluctuations to grow macroscopically before the asymptotic restoring forces prevail.
It is crucial to distinguish between non-normality and reactivity. While $\mathcal{R}>0$ is a sufficient condition for transient growth, non-normality is a necessary but not sufficient condition for reactivity. A system can be non-normal (non-orthogonal eigenvectors) yet non-reactive ($\mathcal{R}<0$), in which case perturbations decay monotonically but potentially with a relaxation time distinct from the asymptotic inverse gap [30, 12]. Thus, $\mathcal{R}$ specifically measures the contribution of the non-symmetric part of the interactions to the local instability.
III.3.1 Cyclic couplings
In this case, the numerical abscissa can be evaluated analytically. In particular, we find that for the relevant portion of parameter space $\gamma_{\mu},\gamma_{\nu}>0$:
@@ III.3.1 Cyclic couplings
Results are summarized in Fig.3(a) and (b), where we present the level curves of $m$ for several values of $\gamma_{\ell}$ and for $c=0,\epsilon$. Notably, Fig.3(b) shows the behavior of the feedforward system as a function of $m$, which qualitatively matches that of the cyclic case, as shown in Fig.3(a). This observation suggests that Eq. (23) may serve as a qualitative reference even for intermediate configurations with $0<c<\epsilon$.
@@
Interestingly, it turns out that reactivity is governed by two distinct quantities: $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, with the absolute value of their ratio being the non-normality parameter. This implies that, although non-normality is needed to trigger reactivity, a fixed non-normality level does not correspond to a single level of reactivity. In other words, the ratio of $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}/\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$ can remain constant while adjusting both numerator and denominator to enhance $m$. This is shown explicitly in Fig.3(a) and (b), where the lines representing constant values of $\xi$ or $\zeta$ cross the level curves of $m$, i.e., while moving along a line with constant non-normality, multiple values of reactivity can be achieved.
@@
Moreover, from Eq. (23) we can define $\Gamma\equiv\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, i.e. the sum of all coupling constants. In what follows we fix the non-normality and distinguish between two regimes: one with weak interactions and weak reactivity ($\Gamma,m\ll 1$), which is illustrated in Fig.3(c) and (e) and one with strong interactions and strong reactivity ($\Gamma,m\gg 1$), which is illustrated in Fig.3 (d) and (f).
@@
Fig.3 illustrates the time-evolution of the total activity $\rho(t)$ $\sqrt{\sum_{i}(x_{i}^{2}+y^{2}_{i})}$, starting from an initial condition close to the bisector, where Eq. (23) holds. In particular, we have chosen parameters allowing the inner systems to be all equivalent, specifically keeping constant the non-normality parameter of the inner systems, $\zeta$. Note the comparison of the evolution of the norm $\rho(t)$ when the reactivity is small (see Fig.3(c)), to a case in which reactivity is large (see Fig.3(d)). As expected, we observe that for large reactivity, the system moves away from the fixed point with vanishing activity before relaxing. From our results, we can conclude that the cyclic system seems to be more reactive than the decoupled system, and that reactivity increases further as $c$ approaches $\epsilon$.
We must note, however, that the reactivity may appear larger for the coupled systems, simply because the coupling itself increases both the reactivity and the non-normality of the system. To isolate the study of the reactivity, we have performed numerical experiments fixing the level of non-normality of the system. In particular, we have chosen parameters in the cyclic system such that $\xi=\zeta$, i.e., we fix $\gamma_{\mu}$ to compensate the effects of $\gamma_{\ell}$. Therefore, Fig.3(e) reports the results for low values of reactivity, and Fig.3(f) is for high values of reactivity, showing larger excursions away from the fixed point. We want to emphasize the two main consequences of our analyses: (i) we have verified that the cyclic system can be reduced to the decoupled case merely through coupling parameter shifts (thus, in the cyclic case, non-reciprocity does not create a new effective dynamical degree of freedom beyond the local E/I module) and (ii) we have shown that the feedforward system undergoes a genuine increase in reactivity when internal non-normality is constant.
III.3.2 Feedforward couplings
In contrast, for the feedforward system, there is no unique global parameter to quantify the total non-normality, and simply summing the parameters controlling the inner and outer non-normality would be arbitrary. While a specific choice of couplings could, in principle, make the reactivity coincide, identifying it would require careful fine-tuning, outside of the scope of the present work. Importantly, the presence of an external non-normal component still enhances the system’s reactive response. In conclusion, the feedforward network actually increases reactivity compared to the decoupled system. In contrast, in the cyclic system, the enhancement of reactivity can be fully accounted for by matching the level of non-normality. Our results show that feedforward architectures provide a robust way to enhance reactivity without increasing local non-normality, a feature that cannot be achieved in symmetric or cyclic networks.
@@ III.4 Analysis of the stochastic evolution of the dynamics
As previously shown in [5, 12], the microscopic underlying system shows very strong finite-size effects. In particular it was shown in [12] using Eq. (3), that the demographic noise induces a minimum of the effective potential is the origin, thus generating a bistability in the system. In presence of strong non-normality, the shear flow along the diagonal, associated with the two nearly-collapsing eigenvectors, facilitates the jumps between the two fixed points and the dynamics resembles the avalanching behaviors observed in neural system, where the global activity undergoes large irregular fluctuations between a low-activity and a high-activity state.
Here, we want to explore the behavior of the coupled system under noise effects. We then couple three E/I populations, where each obeys the system size expansion as in [5]:
@@
$\frac{dy_{i}}{dt}$ $=-\alpha y_{i}+(1-y_{i})f(s_{i})+\sqrt{\alpha y_{i}+(1-y_{i})f(s_{i})}\eta_{y_{i}}.$ (25)
Also in this case, the introduction of noise induces a new stable equilibrium point in zero for all parameter choices, given that $h$ is kept small. Again, when the deterministic stable point bifurcates from zero, bistability appears. Bistability is not a sufficient ingredient to observe avalanching behavior: a high non-normality is needed [12]. However, for too high values of $\xi$ and $\zeta$ the system tends to be trapped in the active minimum. Since we are interested in how the activity propagates across populations, we used the fraction of active excitatory neurons to visualize the trajectories, Fig. 4. Similar plots could be obtained for inhibitory activities.
@@
Once a sufficiently large non-normality level is set, we can distinguish between two separate behaviours, one where the trajectories are incoherent and one where the system shows dynamical patterns, as a function of the parameter $\gamma_{\ell}$. When $\gamma_{\ell}<<\gamma_{\mu},\gamma_{\nu}$ the system lies in an incoherent phase (Fig. 4 (b) and (d)), while when $\gamma_{\ell}\sim\gamma_{\mu},\gamma_{\nu}$ the behavior is more coherent (Fig. 4 (a) and (c)). When the connectivity among units is large enough and the architecture is cyclic, the activation between units is highly correlated (Fig. 4(a)). Instead, in the feedforward case, when the connectivity among units is large enough, the units are activated sequentially. To fix the ideas, let us take two adjacent populations, where the bottom one projects input to –but does not receive any input from– the top one. When the bottom population transits to the active minimum, it activates also the upper population. The process propagates along the chain, giving rise to the step-like trajectories we observe in Fig. 4(c). A discussion about the effective potential induced by the noise and the external couplings is detailed in Supp. Mat. C.
Figure 4: Comparison of cyclic and feedforward setups in the stochastic regime. Top panels show the cyclic configuration; bottom panels show the feedforward setup. Comparing the left and right columns, the system shifts from an incoherent regime (b, d) to either a coherent phase in the cyclic case (a) or step-like trajectories in the feedforward case (c). Parameters: (a) $\gamma_{\ell}=1.13$, $\xi=0.004$; (b) $\gamma_{\ell}=0.13$, $\xi=0.004$; (c) $\gamma_{\ell}=1.13$, $\zeta=0.005$; (d) $\gamma_{\ell}=0.07$, $\zeta=0.005$.
@@ IV Discussion and conclusions
A central motivation of this work, laid out in the Introduction, is that two conceptually distinct sources of non-reciprocity can shape neural dynamics: local dynamical non-normality within each E/I module, and structural non-normality emerging from directed inter-module connectivity. Our results clarify when these two ingredients merely “stack” in an effectively reducible way, and when their combination produces genuinely new transient and stochastic behaviors.
At the local level, each Wilson-Cowan E/I population is tuned to a strongly non-normal regime (captured by the near-alignment of eigenvectors), so that even linearly stable fixed points can display pronounced transient amplification and strong sensitivity to fluctuations. This “inner” mechanism, quantified by the parameter $\zeta$, sets the baseline susceptibility of each unit to perturbations and noise-driven excursions between inactive and active states.
The motif or network level (structural) contribution depends crucially on architecture. In the cyclic (feedback) motif, we find that the effect of coupling can be largely reabsorbed as an effective shift of the excitatory coupling (i.e., a renormalization of parameters), so that the deterministic transition and the core local non-normal mechanism are essentially those of the decoupled system in disguise. In the language of the Introduction, this is a case where directed interactions do not create an independent “structural” non-normal mechanism that fundamentally reshapes the dynamics; instead, the dominant effect is to move the operating point of the same underlying local non-normal module.
By contrast, in the purely feedforward motif, directionality cannot be absorbed into a single effective parameter: the network introduces an additional, explicitly “outer” non-normal component (controlled by $\gamma_{\ell}$) on top of the inner E/I non-normality $\zeta$. This distinction becomes especially sharp in the analysis of reactivity. While non-normality is a prerequisite for reactivity, we show that matching non-normality levels does not uniquely fix reactive amplification; importantly, only the feedforward architecture exhibits a genuine increase in reactivity when inner non-normality is held constant, consistent with the idea that structural non-normality can add an independent amplification channel. In other words, this is precisely the kind of “interplay across scales” anticipated in the Introduction: embedding locally non-normal subsystems into a directed architecture can enhance transient responses in a way that is not captured by local properties alone.
These architectural differences are even more pronounced in the stochastic regime, where demographic noise can generate bistability and, with sufficiently strong non-normality, intermittent avalanche-like excursions. Here, structure governs not only the amount of amplification, but also how activity is distributed in time across populations. With strong coupling and feedback (the cyclic motif), noise-driven transitions are typically coherent, producing synchronized activation. In the feedforward motif, strong coupling instead yields sequential, step-like propagation along the chain, reflecting directional recruitment rather than global coherence. Overall, our results distinguish two roles of non-reciprocity: local non-normality permits large noise-driven excursions, while structural asymmetry determines whether those excursions spread synchronously (feedback) or progressively (feedforward).
Our results also connect to classic work on large random interaction matrices in ecology and neuroscience [26, 33, 7]. In those settings, departing from symmetric (or nearly symmetric) couplings tends to increase eigenvector non-orthogonality and boost transient amplification, making the dynamics much more sensitive to small perturbations. Intuitively, directed (non-reciprocal) interactions can route activity along preferred directions: growth in one subspace can feed into others before relaxation sets in, sustaining irregular fluctuations and, in high dimensions, helping trigger chaotic or strongly time-dependent regimes. From this viewpoint, the “outer” non-normality introduced by directed architecture plays a role analogous to asymmetry in random connectivity: it creates extra pathways for redistributing locally amplified perturbations, hindering simple equilibration and promoting persistent, noise-like dynamics (see also [24, 25]).
Finally, we note limitations and natural extensions. We chose $n=3$ to allow complex eigenvalues and to probe noise-driven time dependence, but we did not observe sustained oscillations; larger $n$ may be needed to obtain more robust quasi-cyclic dynamics. More broadly, an important next step is to move beyond minimal motifs to richer architectures with many more nodes and heterogeneous connectivity, including mixtures of feedforward pathways, recurrent loops, and multiscale feedback, where local and structural non-normal effects could interact in new ways. Testing time-delayed feedback and relaxing the assumption of purely excitatory long-range coupling would further clarify how broadly the separation between inner (local) and outer (structural) non-normality holds in more realistic, large-scale circuits.
Overall, the present framework supports the Introduction’s central claim: non-reciprocity is layered, and understanding neural dynamics requires disentangling when asymmetry acts through local non-normal mechanisms, when it acts through structural non-normal mechanisms, and when their combination yields new transient and noise-driven collective phenomena. We hope this work will encourage further research along these conceptually rich and challenging directions.
Acknowledgments: We acknowledge the Spanish Ministry and Agencia Estatal de Investigación (AEI), MICIN/AEI/10.13039/501100011033, for financial support, Project PID2023-149174NB-I00 funded also by ERDF/EU.